# VL-LN Bench: Towards Long-horizon Goal-oriented Navigation with Active Dialogs

### v3

@@ IV-C Additional Insights
Reliability of evaluation and Oracle The benchmark provides a reliable testbed for IIGN, as humans achieve high success rates. The oracle is also dependable, with the Human–Oracle setting attaining performance comparable to Human–Human (91% vs. 93%), though at the cost of more dialog turns (9.72 vs. 2.04 on average). This discrepancy arises for two reasons: some queries fall outside the oracle’s knowledge or response schema, and users interacting with AI systems often seek additional confirmation (e.g., asking a disambiguation question even after the target has been identified). Although not flawless, the oracle is sufficient for evaluating IIGN, as performance in the Agent–Human and Agent–Oracle settings is similar (16% vs. 17%).
Why do humans fail? As shown in Fig. 4, two factors dominate in the Human-Human setting: (i) Referential ambiguity, where the expression (e.g., “the stairs”) does not uniquely identify the intended instance when multiple same-category objects are visible, causing mismatched grounding; and (ii) Partial observability, where the target is partially occluded so participant (Navigator) assumes a single candidate and commit without disambiguation, leading to a wrong choice. In the Human-Oracle setting, we observe an additional failure mode, Exploration failure, where the human (Navigator) fails to complete the task before the step limit is reached. This typically occurs when the target is difficult to find or the scene is large and complex.

### v4

@@ IV-C Additional Insights
Reliability of evaluation and NPC The benchmark provides a reliable testbed for IIGN, as humans achieve high success rates. The GPT implemented NPC is also dependable, with the Human–GPT setting attaining performance comparable to Human–Human (91% vs. 93%), though at the cost of more dialog turns (9.72 vs. 2.04 on average). This discrepancy arises for two reasons: some queries fall outside the GPT’s knowledge or response schema, and users interacting with AI systems often seek additional confirmation (e.g., asking a disambiguation question even after the target has been identified). Although not flawless, the GPT implemented NPC is sufficient for evaluating IIGN, as performance in the VLLN-D – Human and VLLN-D – GPT settings is similar (16% vs. 17%).
Why do humans fail? As shown in Fig. 4, two factors dominate in the Human-Human setting: (i) Referential ambiguity, where the expression (e.g., “the stairs”) does not uniquely identify the intended instance when multiple same-category objects are visible, causing mismatched grounding; and (ii) Partial observability, where the target is partially occluded so participant (Navigator) assumes a single candidate and commit without disambiguation, leading to a wrong choice. In the Human-GPT setting, we observe an additional failure mode, Exploration failure, where the human (Navigator) fails to complete the task before the step limit is reached. This typically occurs when the target is difficult to find or the scene is large and complex.